# Learning Time-Varying Multi-Region Communications via Scalable Markovian Gaussian Processes

### v2

@@ Appendix D Gaussian Process Regression
$\bullet$ Spectral Mixture (SM) Wilson & Adams (2013): Single-output with
@@
$\bullet$ Multi-Output Squared Exponential (MOSE) Gokcen et al. (2022): Multi-output with $\mathbf{K}_{ij}(t,t^{\prime})=\sigma_{ij}^{2}\exp\left(-\frac{(t-t^{\prime}+\delta_{ij})^{2}}{2l_{ij}^{2}}\right)$.
$\bullet$ Multi-Output Spectral Mixture (MOSM) Parra & Tobar (2017): Multi-output with
@@
$\bullet$ Cross-Spectral Mixture (CSM) Ulrich et al. (2015): Multi-output with
@@ Appendix E Additional Across- and Within-Region Latent Variables
Figure 5 presents the within-region neural activity for both synthetic data and V1-V2 neural spike trains. Figure 6 displays the across- and within-region communications, neural activity, and the estimated temporal delays for the remaining two groups.
Figure 5: Within-region neural activity for both synthetic data and V1-V2 neural spike trains.
Figure 6: The across- and within-region communications, neural activity (ten trials are shown), and the estimated temporal delays for the remaining two groups.

### v4

@@ Appendix D Gaussian Process Regression
$\bullet$ Spectral Mixture (SM) (Wilson & Adams, 2013): Single-output with
@@
$\bullet$ Multi-Output Squared Exponential (MOSE) (Gokcen et al., 2022): Multi-output with $\mathbf{K}_{ij}(t,t^{\prime})=\sigma_{ij}^{2}\exp\left(-\frac{(t-t^{\prime}+\delta_{ij})^{2}}{2l_{ij}^{2}}\right)$.
$\bullet$ Multi-Output Spectral Mixture (MOSM) (Parra & Tobar, 2017): Multi-output with
@@
$\bullet$ Cross-Spectral Mixture (CSM) (Ulrich et al., 2015): Multi-output with
@@ Appendix E Additional Across- and Within-Region Latent Variables
Figure 5 presents the within-region neural activity for both synthetic data and V1-V2 neural spike trains.
Figure 5: (A) Within-region neural activity for synthetic data. (B) Within-region neural activity for V1-V2 neural spike trains (ten trials are shown). (C) The across- and within-region latent variables for the five-region dataset (ten trials are shown).
@@ Appendix F Factor Analysis for Five Regions Spike Trains
Appendix G More Synthetic Data Experiments
Figure 7: Estimated Pairwise Temporal Delays for Synthetic Data with Five Regions. Dashed lines indicate the ground truth delays, red lines show the estimated delays, and shaded areas represent the variance across different runs. The results demonstrate that the model accurately recovers the true delays. The increased MSE observed in Figure 8(A) is attributed to amplitude variability, which is not a practical concern, as the temporal delay patterns are well preserved.
Figure 8: Additional Results for Synthetic Dataset. (A) Model evaluation using MSE and Pearson’s correlation coefficient (CC) between estimated and ground truth delays across different numbers of regions. While MSE increases due to greater amplitude variability with more regions, CC remains stable, indicating reliable recovery of temporal delay patterns. See also Figure 7 for a visualization of estimated delays with five regions. (B) Model evaluation across varying numbers of latent variables, showing stable MSE and CC performance. (C) Model evaluation under different data lengths. Longer sequences yield lower MSE, likely due to improved estimation of underlying delays with more data. (D) Test log-likelihood comparison when the number of latent variables is under-specified, correctly specified, or over-specified. Both under- and over-specification result in lower log-likelihood and higher variance.
Figure 9: Test Log-Likelihood Comparison of ADM, DLAG, MRM-GP, and mp-srSLDS on the Synthetic Dataset used in Section 4.1. ADM consistently outperforms other methods by effectively capturing continuously time-varying temporal delays.
Appendix H Additional Results for Decoding Visual Stimuli in the Five-Region Brain Dataset
Figure 10: Decoding Visual Stimulus Orientation from the Visual Rostrolateral Area (VISrl). We analyze neural data from five regions and extract the learned latent variables. A linear decoder is then used to classify the orientation of visual stimuli ($0^{\circ}$, $90^{\circ}$, and $135^{\circ}$) presented to the mouse during data collection. We evaluate decoding performance using three inputs: (1) raw observed neural activity from Visual Rostrolateral Area, (2) across-region latent variables (representing the communication subspace) of Visual Rostrolateral Area, and (3) within-region latent variables of Visual Rostrolateral Area. The results show that decoding directly from the observed neural data yields the highest test classification accuracy. Among the latent spaces, the communication subspace achieves higher accuracy than the within-region subspace. Notably, both the observed data and the communication subspace perform above random guessing, indicating that orientation information is preserved in the communication subspace of the Visual Rostrolateral Area, which is a region known to be involved in motion and spatial processing. Finally, the drop in accuracy from the observed data to the communication subspace is likely due to information loss from dimensionality reduction.
Appendix I Related Work on Neural Alignment and Task-Dependent Variability Modeling in Multi-Region Anslysis Interaction
Multi-region neural analyses increasingly rely on tools that can jointly characterise activity patterns within each area and the structured co-variability that links them. Williams et al. 2021 gives a geometric foundation for such comparisons by turning representational-similarity heuristics into proper metric spaces; they demonstrate that these “shape metrics” scale to surveys of 48 distinct visual areas in the Allen Brain Observatory, enabling distance-based clustering and embedding of whole-brain activity patterns. Safaie et al. 2023 pushes the multi-region perspective across species, showing that low-dimensional trajectories extracted from motor cortex recordings in both monkeys and mice align in a common latent space and remain predictive of movement even in the absence of overt behaviour, pointing to conserved circuit-level dynamics that transcend individual brains. Complementing these observational studies, Balzani et al., 2022 introduces TAME-GP, a probabilistic manifold model that factorises population variability into within-region and across-region components, imposes Gaussian-process priors for temporal smoothing, and infers single-trial latent dynamics that reveal task-dependent communication subspaces among multiple brain areas